# Multi-user QKD using quotient graph states derived from continuous-variable dual-rail cluster states

### v1

@@ 3.3 Finite Size Analysis
The order in which error correction and parameter estimation are performed affects the finite-size key rate expression [27]. When error correction is performed after parameter estimation, some transmitted signals are disclosed to estimate the channel parameters. Let $N$ be the total number of signals exchanged, and $n$ be the number of signals used for key generation after discarding $N-n$ signals for parameter estimation. The finite-size key rate $K$ is given by:
@@
Alternatively, when error correction is performed before parameter estimation, all measurements can be used for both parameter estimation and key generation, allowing for better estimates of the channel parameters since no data is discarded. In this case, the finite-size key rate $K$ is:
@@
In both expressions, $K_{\infty}(t^{\text{low}},\epsilon_{\text{up}})$ is the asymptotic key rate calculated using conservative estimates of the channel transmission $t$ and excess noise $V_{\epsilon}=t\epsilon$ to account for statistical fluctuations due to the finite sample size. These adjusted parameters are:
$\tau^{\text{low}}$ $=\tau-6.5\,\sigma_{t}$ (10)
@@
where $\sigma_{t}$ and $\sigma_{\epsilon}$ are the standard deviations of the estimators of $t$ and $V_{\epsilon}$, respectively (see Appendix C). The factor $6.5$ corresponds to a confidence level ensuring an error probability of $10^{-10}$. The finite-size correction term $\Delta(n)$ [28]accounts for the statistical fluctuations in the mutual information:
@@
4 Generation of Tripartite States
@@
We consider two configurations: the symmetric GHZ-like state and the downstream access network (DAN). The GHZ-like state is of interest because it has maximum correlation equally distributed among all modes, The DAN configuration is appealing because it can be easily implemented by substituting the two-mode squeezed vacuum with either modulated squeezed or coherent states, and it offers an advantage in producing bipartite keys.
@@
$V_{A}=V_{B}=V_{C}=\text{Diag}\left[\frac{1+2V^{2}}{3V},\ \frac{2+V^{2}}{3V}\right],$
@@
$C_{AB}=-C_{AC}=C_{BC}=\text{Diag}\left[\frac{V^{2}-1}{3V},\ \frac{1-V^{2}}{3V}\right].$
@@
$V_{A}=\text{Diag}\left[\frac{1+V^{2}}{2V},\ \frac{1+V^{2}}{2V}\right],\quad V_{B}=V_{C}=\text{Diag}\left[\frac{(1+V)^{2}}{4V},\ \frac{(1+V)^{2}}{4V}\right],$
$C_{AB}=-C_{AC}=\text{Diag}\left[\frac{V^{2}-1}{2\sqrt{2}V},\ \frac{1-V^{2}}{2\sqrt{2}V}\right],\quad C_{BC}=\text{Diag}\left[-\frac{(V-1)^{2}}{4V},\ -\frac{(1-V)^{2}}{4V}\right].$
@@
4.0.1 Effect of Impurity in Squeezed States
@@
$V_{1}=\text{Diag}\left[V,\ V_{N}+\dfrac{1}{V}\right],\quad V_{2}=V_{3}=\text{Diag}\left[V_{N}+\dfrac{1}{V},+V\right].$
@@
$V_{1}=\text{Diag}\left[V,\ V_{N}+\dfrac{1}{V}\right],\quad V_{2}=\text{Diag}\left[V_{N}+\dfrac{1}{V},V\right],\quad V_{3}=\text{Diag}\left[1,\ 1\right].$
@@
$V_{A}=V_{B}=V_{C}=\text{Diag}\left[\frac{1+2V^{2}+VV_{N}}{3V},\ \frac{2+V^{2}+2VV_{N}}{3V}\right],$
@@
$C_{AB}=-C_{AC}=C_{BC}=\text{Diag}\left[\frac{V^{2}-1-VV_{N}}{3V},\ \frac{1-V^{2}+VV_{N}}{3V}\right].$
@@
$V_{A}=\text{Diag}\left[\frac{1+V^{2}+VV_{N}}{2V},\ \frac{1+V^{2}+VV_{N}}{2V}\right],$
$\quad V_{B}=V_{C}=\text{Diag}\left[\frac{(1+V)^{2}+VV_{N}}{4V},\ \frac{(1+V)^{2}+VV_{N}}{4V}\right],$
$C_{AB}=-C_{AC}=\text{Diag}\left[\frac{V^{2}-1-VV_{N}}{2\sqrt{2}V},\ \frac{1-V^{2}+VV_{N}}{2\sqrt{2}V}\right],$
$\quad C_{BC}=\text{Diag}\left[-\frac{(V-1)^{2}+VV_{N}}{4V},\ -\frac{(1-V)^{2}+VV_{N}}{4V}\right].$
@@
To address this, we introduce an additional mode for each squeezed state to account for the noise, ensuring that the two modes together form a pure bipartite state. Specifically, we consider the impure squeezed state as one of the outputs of a beam splitter fed by two pure squeezed states with x-quadrature variances $\Gamma_{1}$ and $\Gamma_{2}$ at the input. The impure squeezed state are then related to the variances of the input pure squeezed states as follows:
$\Gamma_{1}=V-\frac{V^{3}V_{N}}{\sqrt{V^{3}V_{N}(1+V^{3}V_{N})}},\quad\Gamma_{2}=V+\frac{V^{3}V_{N}}{\sqrt{V^{3}V_{N}(1+V^{3}V_{N})}}.$ (14)
5 Dual-rail cluster state for QKD
In this analysis, we propose a distribution strategy for the modes of a dual-rail cluster state in a multi-node QKD network. Unlike a finite graph defined by a discrete set of nodes, the dual-rail cluster state represents an open-ended, repeating lattice of interconnected modes. Each mode is typically linked to four others, creating an extensive and complex structure that is difficult to handle as is.
@@ 5 Dual-rail cluster state for QKD
The optimal homomorphic graph corresponds to a six-mode graph state, achievable through a coloring strategy detailed in Fig. 2 . The covariance matrix of this six-mode graph state is (refer to Sec: A for derivation):
$\begin{pmatrix}\boldsymbol{V}&\boldsymbol{0}&\boldsymbol{C}&-\boldsymbol{C}&\boldsymbol{C}&\boldsymbol{C}\\ \boldsymbol{0}&\boldsymbol{V}&\boldsymbol{C}&-\boldsymbol{C}&-\boldsymbol{C}&-\boldsymbol{C}\\ \boldsymbol{C}&\boldsymbol{C}&\boldsymbol{V}&\boldsymbol{0}&\boldsymbol{C}&-\boldsymbol{C}\\ -\boldsymbol{C}&-\boldsymbol{C}&\boldsymbol{0}&\boldsymbol{V}&\boldsymbol{C}&-\boldsymbol{C}\\ \boldsymbol{C}&-\boldsymbol{C}&\boldsymbol{C}&\boldsymbol{C}&\boldsymbol{V}&\boldsymbol{0}\\ \boldsymbol{C}&-\boldsymbol{C}&-\boldsymbol{C}&-\boldsymbol{C}&\boldsymbol{0}&\boldsymbol{V}\end{pmatrix}$ (15)
where the diagonal covariance matrix $\boldsymbol{V}$ and the off-diagonal covariance matrix $\boldsymbol{C}$ are defined as follows for pure squeezed states:
$\boldsymbol{V}=\begin{pmatrix}\frac{1+V^{2}}{2V}&0\\ 0&\frac{1+V^{2}}{2V}\end{pmatrix},\quad\boldsymbol{C}=\begin{pmatrix}\frac{1-V^{2}}{4V}&0\\ 0&\frac{V^{2}-1}{4V}\end{pmatrix}.$ (16)
@@
$\boldsymbol{V}=\begin{pmatrix}\frac{1+V^{2}+V_{N}}{2V}&0\\ 0&\frac{1+V^{2}+V_{N}}{2V}\end{pmatrix},\quad\boldsymbol{C}=\begin{pmatrix}\frac{1-V^{2}+VV_{N}}{4V}&0\\ 0&\frac{V^{2}-1-VV_{N}}{4V}\end{pmatrix}.$ (17)
@@
Figure 2: (a) Shows the coloring scheme of dual rail cluster state to obtain the six mode graph state (c). The edges in the graph illustrates the correlations between modes, with edges in two colors indicating the sign of the correlation. Edges of one color represent positive correlations, while edges of the other color represent negative correlations.
5.1 Distribution strategies for six-mode graph state
We identify four primary strategies for distributing quantum states among users, depending on whether the dealer also serves as one of the users or acts solely as a distributor.
@@ 5.1 Distribution strategies for six-mode graph state
    Distribution 1 (Fig.3(a)): The dealer sends mode $b_{1}$ and mode $c_{1}$ to two distant users through a channel characterized by channel transmission $t$ and channel excess noise $\epsilon$ transforming the modes to $\tilde{b}_{1}$ and $\tilde{c}_{1}$, with covariance matrix $\tilde{\boldsymbol{V}}=t(\boldsymbol{V}+\epsilon\mathbb{I}-\mathbb{I})+\mathbb{I}$, $\mathbb{I}$ here is 2X2 identity matrix. The dealer retains the rest of the modes, $a_{1}$,$a_{2}$, $b_{2}$, and $c_{2}$.
@@
Figure 3: The graphs depict various methods for distributing a six-mode graph state among three users. Dashed lines indicate that the correlations (i.e., covariance between the modes $\boldsymbol{C}$) have been reduced by a factor of $\sqrt{t}$, while dotted lines indicate a reduction by a factor of $t$.
6 Quantum conference key distribution protocols
Quantum conference key distribution protocols like standard two user QKD protocols can be broadly categorized into three categories, based on the role of the dealer and the method of key reconciliation. Understanding these protocols is essential for evaluating the security and efficiency of multi-user quantum communication systems. In our analysis, we assume that the channel parameters are symmetric, meaning that all users have identical channel conditions.
@@ 6 Quantum conference key distribution protocols
For the GHZ state described in Sec.4, the covariance matrix given in Eq.13 transforms to:
@@
The key rate for this approach is given by $k^{A}(A:\tilde{B}:\tilde{C})$ (refer to Eq. 3), where the notation $\tilde{i}$ indicates that mode $i$ has been transmitted through a quantum channel.
We explore two distribution strategies for a six-mode graph state intended for three-user conference key generation via direct reconciliation, as illustrated in Figs. 3(a) and 3(b).
@@
From Fig.4(a), we observe that under finite squeezing, the GHZ state is not the optimal multipartite resource for maximizing distance in conference key generation with direct reconciliation under ideal conditions. Among the configurations tested, Distribution 1 achieves the greatest distances, particularly under high excess channel noise ($0.50.08<\epsilon<$ 0.5 SNU). Although trusted noise does not hinder the performance of Distribution 1 or GHZ-based protocols in the asymptotic regime (and may even assist them), it offers no similar advantage to Distribution 2, yet it does not significantly degrade its performance either.
In the finite-size regime, for $n=10^{7}$ measurements and imperfect reconciliation ($\beta=0.95$), Distribution 2 (Fig.3(b)) with key rate $k^{a_{1}}(a_{1}:\tilde{b}_{1}\tilde{b}_{2}:\tilde{c}_{1}\tilde{c}_{2}\,|\,a_{2})$ enables key generation over greater distances for channel noise $0<\epsilon<0.27SNU$; see Fig. 5(a). This is in contrast of what we observe in asymptotic regime where Distribution 1 enables key generation for the largest distances. This is due to the fact that Distribution 2 enables the users and the dealer to use four correlations as opposed to three for Distribution 1 to enable better estimation of the channel parameters (see Sec. C).
@@
The conference key rate for tripartite states under reverse reconciliation is given by $k^{\tilde{B}}(A:\tilde{B}:\tilde{C})$.
@@
The conference key rate for tripartite states in the entanglement-in-the-middle protocol is given by $k^{\tilde{A}}(\tilde{A}:\tilde{B}:\tilde{C})$.
@@
Figure 4: These plots illustrate the maximum tolerable channel losses and noise in asymptotic regime, key generation is possible only for channel parameters lying below the plotted curves. The blue curve corresponds to Distribution 2, the red to Distribution 1, the green to Distribution 3, the brown to Distribution 4, and the black to the GHZ state. Dashed lines represent pure squeezed states, while solid lines indicate impure but trusted squeezed states with $\Delta V=10$. The reconciliation efficiency is $\beta=1$, and the variance of the squeezed quadrature is $V=0.1$ SNU.
Figure 5: These plots illustrate the maximum tolerable channel losses and noise in finite size regime with $n=10^{7}$ measurements, key generation is possible only for channel parameters lying below the plotted curves. The blue curve corresponds to Distribution 2, the red to Distribution 1, the green to Distribution 3, and the black to the GHZ state. With trusted AS noise $\Delta V=10SNU$. The reconciliation efficiency is $\beta=0.95$, and the variance of the squeezed quadrature is $V=0.1SNU$.
@@
Scenario Tripartite-States Key Rate Distribution Key Rate Bipartite key rate post CKA with DR $k^{\tilde{B}}(\tilde{B}:\tilde{C}\,|\,A)$ Distribution 1 $k^{\tilde{b}_{1}}(\tilde{b}_{1}:\tilde{c}_{1}\,|\,a_{1}a_{2})$ Distribution 2 $k^{\tilde{b}_{1}\tilde{b}_{2}}(\tilde{b}_{1}\tilde{b}_{2}:\allowbreak\tilde{c}_{1}\tilde{c}_{2}\,|\,a_{1}a_{2})$ Bipartite key rate with DR post CKA with RR $k^{A}(A:\tilde{C}\,|\,\tilde{B})$ Distribution 1 $k^{a_{1}}(a_{1}:\tilde{c}_{1}\,|\,b_{2}a_{2}c_{2}\tilde{b}_{1})$ Distribution 2 $k^{a_{1}a_{2}}(a_{1}a_{2}:\allowbreak\tilde{c}_{1}\tilde{c}_{2}\,|\,\tilde{b}_{1}\tilde{b}_{2})$ Bipartite key rate with RR post CKA with RR $k^{\tilde{C}}(A:\tilde{C}\,|\,\tilde{B})$ Distribution 1 $k^{\tilde{c}_{1}}(a_{1}:\tilde{c}_{1}\,|\,b_{2}a_{2}c_{2}\tilde{b}_{1})$ Distribution 2 $k^{\tilde{c}_{1}\tilde{c}_{2}}(a_{1}a_{2}:\allowbreak\tilde{c}_{1}\tilde{c}_{2}\,|\,\tilde{b}_{1}\tilde{b}_{2})$ Bipartite key rate post CKA with Mid $k^{\tilde{B}}(\tilde{B}:\tilde{C}\,|\,\tilde{A})$ Distribution 3 $k^{\tilde{b}_{1}}(\tilde{b}_{1}:\tilde{c}_{1}\,|\,a_{2}b_{2}c_{2}\tilde{a}_{1})$ Distribution 4 $k^{\tilde{b}_{1}\tilde{b}_{2}}(\tilde{b}_{1}\tilde{b}_{2}:\allowbreak\tilde{c}_{1}\tilde{c}_{2}\,|\,\tilde{a}_{1}\tilde{a}_{2})$
@@
Figure 6: Maximum tolerable excess noise $\epsilon$ (in SNU) for varying transmission of the channel $t$ for generation of bipartite keys post CKA generation. The blue curve corresponds to Distribution 2, the red to Distribution 1, the green to Distribution 3, the brown to Distribution 4, and the black to the GHZ state. Dashed lines corresponds to finite size keys, while solid lines corresponds to asymptotic keys. For AS noise of $\Delta V=10$, the reconciliation efficiency is $\beta=0.95$, and the variance of the squeezed quadrature is $V=0.1$ SNU.
When comparing the overall multiuser quantum key distribution (QKD) capabilities (CKA and subsequent bipartite key generation) of multipartite states, the six-mode graph state derived from a dual rail cluster state significantly outshines the GHZ state: see Fig.6. For CKA via DR, Distribution 1 supports CKA and bipartite key generation under a broader range of channel conditions than any other considered multipartite state or strategy. For CKA via RR, Distribution 2 excels beyond all others. For CKA via Mid, Distribution 4 facilitates CKA and bipartite keys under more extensive channel conditions. In the finite-size regime, bipartite keys generated using the GHZ state become infeasible for measurements with fewer than $n=10^{8}$ samples. In contrast, protocols employing dual-rail cluster states enable the generation of bipartite keys after the conference key is established, making them applicable to all the protocols introduced. However, the distribution strategy must be carefully tailored to the specific requirements of the quantum network.
@@
In Sec.3, we discussed the utility of analyzing the sum of independent bipartite keys that can be generated by a central user who has higher correlations with other users. This approach is particularly informative in symmetric cases, where the collective key rate helps establish bounds on the symmetric channel parameters required for independent key generation by all the users with the central user. As outlined in Sec. 2, generating independent keys necessitates one user having a stronger correlation with the rest of the users than between the users themselves.
@@
Figure 7: b) Shows maximum tolerable excess noise $\epsilon$ for varying transmission of the channel $t$ for generating independent bipartite keys between the dealer and the remote users. Brown plots correspond to Distribution 5 (a) and the purple plot corresponds to DAN. Dashed line corresponds to finite size keys ($n=10^{7}$) and solid lines corresponds to asymptotic keys. For AS noise of $\Delta V=10$, the reconciliation efficiency is $\beta=0.95$, and the variance of the squeezed quadrature is $V=0.1$ SNU.

### v2

@@ III.3 Finite Size Analysis
The order in which error correction and parameter estimation are performed affects the finite-size key rate expression [36]. When error correction is performed after parameter estimation, some transmitted signals are disclosed to estimate the channel parameters. Let $N$ be the total number of signals exchanged, and $n$ be the number of signals used for key generation after discarding NnN-n signals for parameter estimation. The finite-size key rate $K$ is given by:
@@
Alternatively, when error correction is performed before parameter estimation, all measurements can be used for both parameter estimation and key generation, allowing for better estimates of the channel parameters since no data is discarded [36, 37]. In this case, the finite-size key rate $K$ is:
@@
In both expressions, $K_{\infty}(t^{\text{low}},V_{\epsilon}^{\text{up}})$ is the asymptotic key rate calculated using conservative estimates of the channel transmission $t$ and excess noise $V_{\epsilon}=t\epsilon$ to account for statistical fluctuations due to the finite sample size. These adjusted parameters are:
$t^{\text{low}}$ $=t-6.5\,\sigma_{t}$ (17)
@@
where $\sigma_{t}$ and $\sigma_{\epsilon}$ are the standard deviations of the estimators of $t$ and $V_{\epsilon}$, respectively (see Appendix D). The factor $6.5$ corresponds to a confidence level ensuring an error probability of $10^{-10}$. The finite-size correction term $\Delta(n)$ [38]accounts for the statistical fluctuations in the mutual information:
@@
IV Generation of Tripartite States
@@
We consider two configurations: the symmetric GHZ-like state and the downstream access network (DAN). The GHZ-like state is of interest because it has maximum correlation equally distributed among all modes. The DAN configuration is appealing because it can be easily implemented by substituting the two-mode squeezed vacuum with either modulated squeezed or coherent states, and it offers an advantage in producing bipartite keys.
@@
$V_{A}=V_{B}=V_{C}=\text{Diag}\left[\frac{1+2V^{2}}{3V},\ \frac{2+V^{2}}{3V}\right],$ (21)
@@
$C_{AB}=-C_{AC}=C_{BC}=\text{Diag}\left[\frac{V^{2}-1}{3V},\ \frac{1-V^{2}}{3V}\right].$ (22)
@@
$V_{A}$ $=\mathrm{Diag}\!\left[\tfrac{1+V^{2}}{2V},\,\tfrac{1+V^{2}}{2V}\right],$ (23)
$V_{B}$ $=V_{C}=\mathrm{Diag}\!\left[\tfrac{(1+V)^{2}}{4V},\,\tfrac{(1+V)^{2}}{4V}\right],$
$C_{AB}$ $=-C_{AC}=\mathrm{Diag}\!\left[\tfrac{V^{2}-1}{2\sqrt{2}V},\,\tfrac{1-V^{2}}{2\sqrt{2}V}\right],$
$C_{BC}$ $=\mathrm{Diag}\!\left[-\tfrac{(V-1)^{2}}{4V},\,-\tfrac{(1-V)^{2}}{4V}\right].$
@@
IV.0.1 Effect of Impurity in Squeezed States
@@
$V_{1}$ $=\mathrm{Diag}\!\left[V,\,V_{N}+\tfrac{1}{V}\right],$ (24)
$V_{2}$ $=V_{3}=\mathrm{Diag}\!\left[V_{N}+\tfrac{1}{V},\,V\right].$
@@
$V_{1}$ $=\mathrm{Diag}\!\left[V,\,V_{N}+\tfrac{1}{V}\right],$ (25)
$V_{2}$ $=\mathrm{Diag}\!\left[V_{N}+\tfrac{1}{V},\,V\right],$
$V_{3}$ $=\mathrm{Diag}\!\left[1,\,1\right].$
@@
$V_{A}$ $=V_{B}=V_{C}=\mathrm{Diag}\!\left[\tfrac{1+2V^{2}+VV_{N}}{3V},\,\tfrac{2+V^{2}+2VV_{N}}{3V}\right],$ (26)
@@
$C_{AB}$ $=-C_{AC}=C_{BC}=\mathrm{Diag}\!\left[\tfrac{V^{2}-1-VV_{N}}{3V},\,\tfrac{1-V^{2}+VV_{N}}{3V}\right].$ (27)
@@
$V_{A}$ $=\mathrm{Diag}\!\left[\tfrac{1+V^{2}+VV_{N}}{2V},\,\tfrac{1+V^{2}+VV_{N}}{2V}\right],$ (28)
$V_{B}$ $=V_{C}=\mathrm{Diag}\!\left[\tfrac{(1+V)^{2}+VV_{N}}{4V},\,\tfrac{(1+V)^{2}+VV_{N}}{4V}\right],$
$C_{AB}$ $=-C_{AC}=\mathrm{Diag}\!\left[\tfrac{V^{2}-1-VV_{N}}{2\sqrt{2}V},\,\tfrac{1-V^{2}+VV_{N}}{2\sqrt{2}V}\right],$
$C_{BC}$ $=\mathrm{Diag}\!\left[-\tfrac{(V-1)^{2}+VV_{N}}{4V},\,-\tfrac{(1-V)^{2}+VV_{N}}{4V}\right].$
@@
To address this, we introduce an additional mode for each squeezed state to account for the noise, ensuring that the two modes together form a pure bipartite state. Specifically, we consider the impure squeezed state as one of the outputs of a beam splitter fed by two pure squeezed states with x-quadrature variances $\Gamma_{1}$ and $\Gamma_{2}$ at the input. The impure squeezed state is then related to the variances of the input pure squeezed states as follows:
$\Gamma_{1}$ $=V-\frac{V^{3}V_{N}}{\sqrt{V^{3}V_{N}\bigl(1+V^{3}V_{N}\bigr)}},$ (29)
$\Gamma_{2}$ $=V+\frac{V^{3}V_{N}}{\sqrt{V^{3}V_{N}\bigl(1+V^{3}V_{N}\bigr)}}.$
V Dual-rail cluster state for QKD
In this analysis, we propose a distribution strategy for the modes of a dual-rail cluster state in a multi-node QKD network. Unlike a finite graph defined by a discrete set of nodes, the dual-rail cluster state represents an open-ended, repeating lattice of interconnected modes. Each mode is typically linked to four others, creating an extensive and complex structure that is difficult to handle as it is.
@@ V Dual-rail cluster state for QKD
The optimal homomorphic graph corresponds to a six-mode graph state, achievable through a coloring strategy detailed in Fig. 4 . The covariance matrix of this six-mode graph state is (refer to Sec. A for derivation):
$\begin{pmatrix}\bm{V}&\bm{0}&\bm{C}&-\bm{C}&\bm{C}&\bm{C}\\ \bm{0}&\bm{V}&\bm{C}&-\bm{C}&-\bm{C}&-\bm{C}\\ \bm{C}&\bm{C}&\bm{V}&\bm{0}&\bm{C}&-\bm{C}\\ -\bm{C}&-\bm{C}&\bm{0}&\bm{V}&\bm{C}&-\bm{C}\\ \bm{C}&-\bm{C}&\bm{C}&\bm{C}&\bm{V}&\bm{0}\\ \bm{C}&-\bm{C}&-\bm{C}&-\bm{C}&\bm{0}&\bm{V}\end{pmatrix}$ (30)
where the diagonal covariance matrix $\bm{V}$ and the off-diagonal covariance matrix $\bm{C}$ are defined as follows for pure squeezed states:
$\bm{V}=\begin{pmatrix}\frac{1+V^{2}}{2V}&0\\ 0&\frac{1+V^{2}}{2V}\end{pmatrix},\quad\bm{C}=\begin{pmatrix}\frac{1-V^{2}}{4V}&0\\ 0&\frac{V^{2}-1}{4V}\end{pmatrix}$. (31)
@@
$\bm{V}$ $=\begin{pmatrix}\tfrac{1+V^{2}+V_{N}}{2V}&0\\[6.0pt] 0&\tfrac{1+V^{2}+V_{N}}{2V}\end{pmatrix}$, (32)
$\bm{C}$ $=\begin{pmatrix}\tfrac{1-V^{2}+VV_{N}}{4V}&0\\[6.0pt] 0&\tfrac{V^{2}-1-VV_{N}}{4V}\end{pmatrix}$.
@@
Figure 4: (a) Shows the coloring scheme of dual rail cluster state to obtain the six mode graph state (b). The edges in the graph illustrates the correlations between modes, with edges in two colors indicating the sign of the correlation. Edges of one color represent positive correlations, while edges of the other color represent negative correlations.
V.1 Distribution strategies for six-mode graph state
In our investigation of six-mode graph states for QKD, we explored a wide array of strategies for distributing the quantum states among users. However, due to space constraints and to focus on the most impactful designs, here we present only the strategies that proved to be the most noteworthy in terms of operational performance and security. We categorize the strategies based on whether the dealer also participates as a user or acts solely as a distributor.
@@ V.1 Distribution strategies for six-mode graph state
    Distribution 1 (Fig. 5(a)): The dealer sends mode $b_{1}$ and mode $c_{1}$ to two distant users through a channel characterized by channel transmission $t$ and channel excess noise $\epsilon$ transforming the modes to $\tilde{b}_{1}$ and $\tilde{c}_{1}$, with covariance matrix $\tilde{\bm{V}}=t(\bm{V}+\epsilon\mathbb{I}-\mathbb{I})+\mathbb{I}$, $\mathbb{I}$ here is $2\times 2$ identity matrix. The dealer retains the rest of the modes, $a_{1}$$a_{2}$, $b_{2}$, and $c_{2}$.
@@
Figure 5: The graphs depict various methods for distributing a six-mode graph state among three users. Dashed lines indicate that the correlations (i.e., covariance between the modes $\bm{C}$) have been reduced by a factor of $\sqrt{t}$, while dotted lines indicate a reduction by a factor of $t$.
VI Quantum conference key distribution protocols
Quantum conference key distribution protocols like standard two-user QKD protocols can be broadly categorized into three categories, based on the role of the dealer and the method of key reconciliation. Understanding these protocols is essential for evaluating the security and efficiency of multi-user quantum communication systems. In our analysis, we assume that the channel parameters are symmetric, meaning that all users have identical channel conditions.
@@ VI Quantum conference key distribution protocols
For the GHZ state described in Sec. IV, the covariance matrix given in Eq. 20 transforms to:
@@
[…]
We explore two distribution strategies for a six-mode graph state intended for three-user conference key generation via direct reconciliation, as illustrated in Figs. 5(a) and 5(b).
@@
From Fig. 6(a), we observe that under finite squeezing, the GHZ state is not the optimal multipartite resource for maximizing distance in conference key generation with direct reconciliation under ideal conditions. Among the configurations tested, Distribution 1 achieves the greatest distances, particularly under high excess channel noise ($0.08<\epsilon<0.5$ SNU). Although trusted noise does not hinder the performance of Distribution 1 or GHZ-based protocols in the asymptotic regime (and may even assist them), it offers no similar advantage to Distribution 2, yet it does not significantly degrade its performance either.
In the finite-size regime, for $n=10^{7}$ measurements and imperfect reconciliation ($\beta=0.95$), Distribution 2 (Fig. 5(b)) with key rate $k^{a_{1}}(a_{1}:\tilde{b}_{1}\tilde{b}_{2}:\tilde{c}_{1}\tilde{c}_{2}\,|\,a_{2})$ enables key generation over greater distances for channel noise $0<\epsilon<0.27SNU$; see Fig. 7(a). This is in contrast to what we observe in the asymptotic regime where Distribution 1 enables key generation for the largest distances. This is due to the fact that Distribution 2 enables the users and the dealer to use four correlations as opposed to three for Distribution 1 to enable better estimation of the channel parameters (see Sec. D).
@@
[…]
@@
[…]
@@
Figure 6: These plots illustrate the maximum tolerable channel losses and noise in asymptotic regime, key generation is possible only for channel parameters lying below the plotted curves. The blue curve corresponds to Distribution 2, the red to Distribution 1, the green to Distribution 3, the brown to Distribution 4, and the black to the GHZ state. Dashed lines represent pure squeezed states, while solid lines indicate impure but trusted squeezed states with $V_{N}=10$. The reconciliation efficiency is $\beta=1$, and the variance of the squeezed quadrature is $V=0.1$ SNU.
Figure 7: These plots illustrate the maximum tolerable channel losses and noise in finite size regime with $n=10^{7}$ measurements, key generation is possible only for channel parameters lying below the plotted curves. The blue curve corresponds to Distribution 2, the red to Distribution 1, the green to Distribution 3, and the black to the GHZ state. With trusted AS noise $V_{N}=10SNU$. The reconciliation efficiency is $\beta=0.95$, and the variance of the squeezed quadrature is $V=0.1SNU$.
@@
Scenario Tripartite-States Key Rate Distribution Key Rate Expression
Bipartite key rate post CKA with DR $k^{\tilde{B}}(\tilde{B}:\tilde{C}\,|\,A)$ Distribution 1 $k^{\tilde{b}_{1}}(\tilde{b}_{1}:\tilde{c}_{1}\,|\,a_{1}a_{2})$
$k^{\tilde{b}_{1}\tilde{b}_{2}}(\tilde{b}_{1}\tilde{b}_{2}:\allowbreak\tilde{c}_{1}\tilde{c}_{2}\,|\,a_{1}a_{2})$
[…]
$k^{a_{1}a_{2}}(a_{1}a_{2}:\allowbreak\tilde{c}_{1}\tilde{c}_{2}\,|\,\tilde{b}_{1}\tilde{b}_{2})$
[…]
$k^{\tilde{c}_{1}\tilde{c}_{2}}(a_{1}a_{2}:\allowbreak\tilde{c}_{1}\tilde{c}_{2}\,|\,\tilde{b}_{1}\tilde{b}_{2})$
[…]
$k^{\tilde{b}_{1}\tilde{b}_{2}}(\tilde{b}_{1}\tilde{b}_{2}:\allowbreak\tilde{c}_{1}\tilde{c}_{2}\,|\,\tilde{a}_{1}\tilde{a}_{2})$
Figure 8: Maximum tolerable excess noise $\epsilon$ (in SNU) versus transmission $t$ for bipartite keys post CKA generation. Blue: Distribution 2; red: Distribution 1; green: Distribution 3; brown: Distribution 4; black: GHZ state. Dashed = finite-size keys, solid = asymptotic. Parameters: $V_{N}=10$, $\beta=0.95$, squeezed variance $V=0.1$ SNU.
When comparing the overall multiuser QKD capabilities (CKA and subsequent bipartite key generation) of multipartite states, the six-mode graph state derived from a dual rail cluster state significantly outshines the GHZ state: see Fig. 8. For CKA via DR, Distribution 1 supports CKA and bipartite key generation under a broader range of channel conditions than any other considered multipartite state or strategy. For CKA via RR, Distribution 2 excels beyond all others. For CKA via Mid, Distribution 4 facilitates CKA and bipartite keys under more extensive channel conditions. In the finite-size regime, bipartite keys generated using the GHZ state become infeasible for measurements with fewer than $n=10^{8}$ samples. In contrast, protocols employing dual-rail cluster states enable the generation of bipartite keys after the conference key is established, making them applicable to all the protocols introduced. However, the distribution strategy must be carefully tailored to the specific requirements of the quantum network.
@@
In Sec. III, we discussed the utility of analyzing the sum of independent bipartite keys that can be generated by a central user who has higher correlations with other users. This approach is particularly informative in symmetric cases, where the collective key rate helps establish bounds on the symmetric channel parameters required for independent key generation by all the users with the central user. As outlined in Sec. II, generating independent keys necessitates one user having a stronger correlation with the rest of the users than between the users themselves.
@@
Figure 9: b) Shows maximum tolerable excess noise $\epsilon$ for varying transmission of the channel $t$ for generating independent bipartite keys between the dealer and the remote users. Orange plots correspond to Distribution 5 (a) and the purple plot corresponds to DAN. Dashed line corresponds to finite size keys ($n=10^{7}$) and solid lines corresponds to asymptotic keys. For AS noise of $V_{N}=10$, the reconciliation efficiency is $\beta=0.95$, and the variance of the squeezed quadrature is $V=0.1$ SNU.